The position of the point of equilateral are respectively (0,0),(a,11) ,(b,37)so what will be the value of a and b?

Grade:10

1 Answers

Aditya Gupta
5 years ago
the distance between all points should be equal.
so, 11^2+a^2=37^2+b^2=(37 – 11)^2+(b-a)^2=26^2+(b-a)^2
solve these equations by isolating a, to obtain
a=21root{3} and b=5root3
or a= – 21root{3} and b= – 5root3